# MyShadeQuoteStyleleftmargin=0pt, rightmargin=0pt, backgroundcolor=gray!25, linewidth=0pt, skipbelow=skipabove=

### v3

@@ Abstract.
ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software reliabilityccs: Software and its engineering Software safetyccs: Software and its engineering Parsersccs: Software and its engineering Software maintenance toolsccs: Security and privacy Vulnerability scannersccs: Security and privacy Software security engineeringccs: Software and its engineering Source code generation
@@ 1. Introduction
Vulnerability detection is a crucial task in software engineering. Recent Deep Learning-based Vulnerability Detection (DLVD) models (Li et al., 2018; Fu and Tantithamthavorn, 2022; Zhou et al., 2019; Chakraborty et al., 2020; Li et al., 2021; Wang et al., 2023b; Lu et al., 2024) have drawn more attention from research and industry communities due to their promising performance. However, these models suffer greatly from 1) a lack of sizable datasets and 2) discrepancies between the distribution of training and testing datasets. Data Sampling techniques (Yang et al., 2023; Ganz et al., 2023) can alleviate the former problem to an extent by upsampling vulnerable samples to balance the dataset and optionally adding more clean samples, which are abundant. However, Data Augmentation presents a more promising solution as it can help with both problems by generating large diverse datasets. For instance, Zhang (Zhang, 2021) and Liu et al. (Liu et al., 2024) focus on generating specific type(s) of bug/vulnerability. Similarly, Nong et al. proposed VulGen (Nong et al., 2023) and VGX (Nong et al., 2024) to mine single-statement vulnerability patterns and inject them into clean samples to generate vulnerable samples. Despite their contributions, these methods have limitations. They either generate only specific types of vulnerabilities (Zhang, 2021; Nong et al., 2024, 2023) or require sizable datasets to learn models for generating vulnerable code or pinpoint the location to inject the vulnerability (Nong et al., 2022, 2024, 2023), making them impractical for adoption in a real-world scenario.
It is worth noting that recent works have also explored vulnerability detection using LLMs (Du et al., 2024; Lu et al., 2024; Ding et al., 2024; Lin and Mohaisen, 2025), which pose as competitors to DLVD models. However, adopting LLM-based vulnerability detection methods poses various challenges and limitations that make DLVD models still highly relevant: 1) Data Privacy Concerns: in the industry, there are privacy concerns regarding the use of LLMs for vulnerability detection and essentially exposing private commercial code to LLMs; 2) Limited Robustness to Domain-specifc Vulnerabilities: LLMs require robust datasets to be used alongside various techniques, thus they usually struggle with project and domain-specifc vulnerabilities not present in the training data; 3) High Computational Costs: LLMs are resource-intensive to run while VulScribeR uses LLMs to augment vulnerable samples once and the trained DLVD models can be deployed while requiring significantly less compute and memory, making them easier to be deployed on-premise or in a constrained environment. Consequently, it is important to augment vulnerable data to improve the performance of DLVD models.
@@
To generate realistic code, we employ Retrieval-Augmented Generation (RAG) in Injection and Extension. More specifically, in Injection, we retrieve vulnerable samples that are similar to the clean sample into which the vulnerable segment will be injected. Additionally, to enhance the diversity of the generated vulnerable samples, we employ a clustering process to ensure the data in all clusters have the choice to be retrieved during RAG. We employ RAG and clustering in Extension similarly.
@@
    Results: RAG component makes a significant contribution to Injection and Extension strategies.
@@ 2.2. Source Code Augmentation with RAG and LLMs
Leveraging LLMs for source code augmentation is getting popular and it has been used for data augmentation for Semantic Code Search (Wang et al., 2023a) and Code Generation (Chen et al., 2024), both of which heavily rely on RAG. Wang et al. (Wang et al., 2023a) leverages RAG to retrieve similar $query-code$ pairs, then for each pair they instruct ChatGPT (OpenAI, 2022) to change the query and code in separate prompts using predefined rules included in the prompts. Then, they filter the augmented pairs using an encoder model that calculates their similarity score, removing the low-scoring pairs. Chen et al. (Chen et al., 2024) utilized Code Search to augment data for Code Generation. Specifically, they retrieve $context-function$ pairs, where the context can be either a function header or a comment, to populate prompt templates that instruct the LLM to generate code based on the given context and retrieved function. Their architecture uses three components, namely Retriever, Formulator, and Generator. Similar to previous studies, we also employ RAG to retrieve relevant context to enhance the code generation, while different from them, we focus on augmenting vulnerabilities and we employ a clustering process to ensure the diversity of generated data.
@@ 3. Methodology
  • Generator In the Generator, we use the instantiated templates to prompt the LLM to generate vulnerable code samples. If the response does not contain a code snippet or we face an API error, a retry mechanism is activated to feed the prompt to the LLM up to 3 times to receive a response that contains a code snippet, otherwise, we move on to the next prompt.
  • Verifier The goal of the Verifier is to have control over the quality of the generated code, as there’s no guarantee that the LLM produces compilable code. We use Joern’s (QwietAI, [n. d.]), a fuzzy C language parser, which is based on Antlr’s (Parr and Quong, 1995) C Parser, to filter out generated codes that contain severe syntax errors. The rationale behind using a fuzzy parser instead of a strict parser is that in data augmentation for various machine learning tasks, the generated data does not need to conform to the original data’s strict standards (DeVries and Taylor, 2017; Zhang et al., 2018; Guo et al., 2019; Li et al., 2022; Dong et al., 2025; Olsson et al., 2021; Yun et al., 2019). Note that our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. As such, completely correct code is not necessary for training models (Wang et al., 2024), and augmented code can improve the performance of the models even when they slightly break the source code syntax (Li et al., 2022; Dong et al., 2025); in fact, data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). We use a parser instead of a compiler as we wish to keep the post-processing lightweight, as compilation needs significant efforts, such as identifying the dependencies and resolving type errors, which are unavailable in the original datasets, as they only contain functions. Nevertheless, we acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e, a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible. As a result, the fuzzy C Parser filters out 2%-13% of the generated data.
@@ 3.1.2. Injection
Following clustering, we create an index on each cluster using Lucene to facilitate efficient search (lines 7-10). For each clean sample, we search for the most similar vulnerable sample within each cluster using the cluster index, and store the resulting clean-vul pairs for further processing (lines 12-17). To find similar vulnerable samples for each clean sample, we use the BM-25 algorithm (Robertson and Zaragoza, 2009), as previous studies have shown that there is no significant difference between sparse methods like BM-25 and dense methods like CodeBERT when leveraging RAG (Chen et al., 2024; Gao et al., 2023; Nashid et al., 2023). We use Lucene’s implementation of BM-25 for this study.
@@
Input: $N$: Number of pairs to sample
@@
Output: $D$: Dataset with $N$ Clean-Vulnerable pairs
@@
[…]
@@
[…]
@@
25       $sort\_by\_score(\textit{clustered\_pairs}[cluster\_id],"desc")$
@@ 33      $\textit{index}\leftarrow i//\textit{G}$
[…]
@@ 4.7.2. Approach of RQ2
In this RQ, we investigate the usefulness of our proposed RAG for the Injection and Extension strategies. Specifically, we eliminate the Retriever component, and instead, we match the clean and vulnerable samples randomly to examine the effect of RAG (i.e. Injection w/o Retriever and Extension w/o Retriever). Moreover, we also conduct an ablation study on the clustering phase of the Retriever component to examine the effect of diversity on the quality of the generated samples. To do so, we remove the clustering phase from the Retriever component and search for similar samples for each clean input item within all the vulnerable items instead of within each cluster. Then we take the top 5 retrieved items to end up with a similar number of $clean-vul$ pairs. In the sampling phase, we sort all the pairs based on their relevance score and start from the top. We denote the Injection strategy without clustering phase as Injection w/o Clustering and the Extension strategy without clustering as Extension w/o Clustering.

### v4

# VulScribeR: Exploring RAG-based Vulnerability Augmentation with LLMs

@@ Abstract.
$c$s: Software and its engineering Software testing and debuggingccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software reliabilityccs: Software and its engineering Software safetyccs: Software and its engineering Parsersccs: Software and its engineering Software maintenance toolsccs: Security and privacy Vulnerability scannersccs: Security and privacy Software security engineeringccs: Software and its engineering Source code generation
@@ 1. Introduction
Vulnerability detection is a crucial task in software engineering. Recent Deep Learning-based Vulnerability Detection (DLVD) models (Li et al., 2018; Fu and Tantithamthavorn, 2022; Zhou et al., 2019; Chakraborty et al., 2020; Li et al., 2021; Wang et al., 2023b; Lu et al., 2024) have drawn more attention from research and industry communities due to their promising performance. However, these models suffer greatly from 1) a lack of sizable datasets and 2) discrepancies between the distribution of training and testing datasets. Data Sampling techniques (Yang et al., 2023; Ganz et al., 2023) can alleviate the former problem to an extent by upsampling vulnerable samples to balance the dataset and optionally adding more clean samples, which are abundant. However, Data Augmentation presents a more promising solution as it can help with both problems by generating large diverse datasets. For instance, Zhang (Zhang, 2021) and Liu et al. (Liu et al., 2024) focus on generating specific type(s) of bug/vulnerability. Similarly, Nong et al. proposed VulGen (Nong et al., 2023) and VGX (Nong et al., 2024) to mine single-statement vulnerability patterns and inject them into clean samples to generate vulnerable samples. Alternatively, Daneshvar et al. (Daneshvar et al., 2025) evaluated mixup-inspired (Zhang et al., 2018) augmentation techniques for token-based DLVD and proposed conditioned variants for augmenting vulnerabilities at the representation level. Despite their contributions, these methods have limitations. They either generate only specific types of vulnerabilities (Zhang, 2021; Nong et al., 2024, 2023) or require sizable datasets to learn models for generating vulnerable code or pinpointing the location to inject the vulnerability (Nong et al., 2022, 2024, 2023), or they are limited to a certain class of DLVD models while yielding minimal performance gains (Daneshvar et al., 2025), making them impractical for adoption in a real-world scenario.
It is worth noting that recent works have also explored vulnerability detection using LLMs (Du et al., 2024; Lu et al., 2024; Ding et al., 2024; Lin and Mohaisen, 2025), which pose as competitors to DLVD models. However, adopting LLM-based vulnerability detection methods poses various challenges and limitations (Abdali et al., 2025; Yao et al., 2024) that make DLVD models still highly relevant: 1) Data Privacy Concerns: in the industry, there are privacy concerns regarding the use of LLMs for vulnerability detection and essentially exposing private commercial code to LLMs; 2) Limited Robustness to Domain-specifc Vulnerabilities: LLMs require robust datasets to be used alongside various techniques, thus they usually struggle with project and domain-specifc vulnerabilities not present in the training data; 3) High Computational Costs: LLMs are resource-intensive to run while VulScribeR uses LLMs to augment vulnerable samples once and the trained DLVD models can be deployed while requiring significantly less compute and memory, making them easier to be deployed on-premise or in a constrained environment. Consequently, it is important to augment vulnerable data to improve the performance of DLVD models.
@@
To generate realistic code, we employ Retrieval-Augmented Generation (RAG) in Injection and Extension. More specifically, in Injection, we retrieve vulnerable samples that are similar to the clean sample into which the vulnerable segment will be injected. Additionally, to enhance the diversity of the generated vulnerable samples, we employ a clustering process to ensure that the data in all clusters have the choice to be retrieved during RAG. We employ RAG and clustering in Extension similarly.
@@
    Results: RAG component makes a significant contribution to Injection and Extension strategies, accounting for 4.99% and 10.77% of their respective improvements.
@@ 2.2. Source Code Augmentation with RAG and LLMs
Leveraging LLMs for source code augmentation is getting popular and it has been used for data augmentation for Semantic Code Search (Wang et al., 2023a) and Code Generation (Chen et al., 2024), both of which heavily rely on RAG. Wang et al. (Wang et al., 2023a) leverages RAG to retrieve similar querycodequery-code pairs, then for each pair they instruct ChatGPT (OpenAI, 2022) to change the query and code in separate prompts using predefined rules included in the prompts. Then, they filter the augmented pairs using an encoder model that calculates their similarity score, removing the low-scoring pairs. Chen et al. (Chen et al., 2024) utilized Code Search to augment data for Code Generation. Specifically, they retrieve contextfunctioncontext-function pairs, where the context can be either a function header or a comment, to populate prompt templates that instruct the LLM to generate code based on the given context and retrieved function. Their architecture uses three components, namely Retriever, Formulator, and Generator. Similar to previous studies, we also employ RAG to retrieve relevant context to enhance the code generation, while different from them, we focus on augmenting vulnerabilities and we employ a clustering process to ensure the diversity of generated data.
@@ 3. Methodology
  • Generator In the Generator, we use the instantiated templates to prompt the LLM to generate vulnerable code samples. If the response does not contain a code snippet or we face an API error, a retry mechanism is activated to feed the prompt to the LLM up to three times to receive a response that contains a code snippet; otherwise, we move on to the next prompt.
  • Verifier The goal of the Verifier [^1] Verifier is a carefully chosen generic term; in this work, it is only a fuzzy parser that filters code with severe syntax errors. is to have control over the quality of the generated code, as there’s no guarantee that the LLM produces compilable code. We use Joern’s (QwietAI, [n. d.]), a fuzzy C language parser, which is based on Antlr’s (Parr and Quong, 1995) C Parser, to filter out generated codes that contain severe syntax errors. The rationale behind using a fuzzy parser instead of a strict parser is that in data augmentation for various machine learning tasks, the generated data does not need to conform to the original data’s strict standards (DeVries and Taylor, 2017; Zhang et al., 2018; Guo et al., 2019; Li et al., 2022; Dong et al., 2025; Olsson et al., 2021; Yun et al., 2019; Daneshvar et al., 2025). Note that our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. As such, completely correct code is not necessary for training models (Wang et al., 2024), and augmented code can improve the performance of the models even when they slightly break the source code syntax (Li et al., 2022; Dong et al., 2025); in fact, data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). We use a parser instead of a compiler as we wish to keep the post-processing lightweight, as compilation needs significant efforts, such as identifying the dependencies and resolving type errors, which are unavailable in the original datasets, as they only contain functions. Nevertheless, we acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e, a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible. As a result, the fuzzy C Parser filters out 2%-13% of the generated data.
@@ 3.1.2. Injection
Following clustering, we create an index on each cluster using Lucene to facilitate efficient search (lines 7-10). For each clean sample, we search for the most similar vulnerable sample within each cluster using the cluster index, and store the resulting clean-vul pairs for further processing (lines 12-17). To find similar vulnerable samples for each clean sample, we use the BM-25 algorithm (Robertson and Zaragoza, 2009), as previous studies have shown that there is no significant difference between sparse methods like BM-25 and dense methods like CodeBERT when leveraging RAG (Chen et al., 2024; Gao et al., 2023; Nashid et al., 2023; He et al., 2025). We use Lucene’s implementation of BM-25 for this study.
@@
Input: NN: Number of pairs to sample
@@
Output: $D$: Dataset with NN Clean-Vulnerable pairs
@@
11 $search\_engine\_index\leftarrow\text{index}(\textit{clustered\_vuls}[\textit{cluster\_id}])$
12 $\textit{indices}[\textit{cluster\_id}].\text{append}(search\_engine\_index)$
@@ 11 $search\_engine\_index\leftarrow\text{index}(\textit{clustered\_vuls}[\textit{cluster\_id}])$
17  for $\textit{cluster\_id}=0$ to G1G-1 do
18     $\textit{vul}\leftarrow\text{search}(\textit{indices}[\textit{cluster\_id}],\textit{clean})$
19    $\textit{clustered\_pairs}[\textit{cluster\_id}].\text{append}((\textit{clean},\textit{vul}))$
@@ 18     $\textit{vul}\leftarrow\text{search}(\textit{indices}[\textit{cluster\_id}],\textit{clean})$
29 $\textit{clustered\_pairs}\leftarrow\text{sort\_arrays\_by\_size}(\textit{clustered\_pairs},\text{"desc"})$
@@ 33 $\textit{index}\leftarrow i//\textit{G}$
34 if $\textit{index}<\text{size}(\textit{pairs})$ then
@@ 4.7.2. Approach of RQ2
In this RQ, we investigate the usefulness of our proposed RAG for the Injection and Extension strategies. Specifically, we eliminate the Retriever component, and instead, we match the clean and vulnerable samples randomly to examine the effect of RAG (i.e. Injection w/o Retriever and Extension w/o Retriever). Moreover, we also conduct an ablation study on the clustering phase of the Retriever component to examine the effect of diversity on the quality of the generated samples. To do so, we remove the clustering phase from the Retriever component and search for similar samples for each clean input item within all the vulnerable items instead of within each cluster. Then we take the top 5 retrieved items to end up with a similar number of cleanvulclean-vul pairs. In the sampling phase, we sort all the pairs based on their relevance score and start from the top. We denote the Injection strategy without clustering phase as Injection w/o Clustering and the Extension strategy without clustering as Extension w/o Clustering.